Randomly selected 20 student cars have ages with a mean of 7.3 years and a standard deviation of 3.4 years, while randomly selected 11 aculty cars have ages with a mean of 5 years and a standard deviation of 3.7 years. b) The test statistic is (c) The p-value is
Answer #1

Data:        

n1 = 20       

n2 = 11       

x1-bar = 7.3       

x2-bar = 5       

s1 = 3.4       

s2 = 3.7       

Hypotheses:        

Ho: μ1 = μ2        

Ha: μ1 ≠ μ2        

Decision Rule:        

α = 0.05       

Degrees of freedom = 20 + 11 - 2 = 29      

Lower Critical t- score = -2.045229611       

Upper Critical t- score = 2.045229611       

Reject Ho if |t| > 2.045229611       

Test Statistic:        

Pooled SD, s = √[{(n1 - 1) s1^2 + (n2 - 1) s2^2} / (n1 + n2 - 2)] =   √(((20 - 1) * 3.4^2 + (11 - 1) * 3.7^2)/(20 + 11 -2)) =     3.506

SE = s * √{(1 /n1) + (1 /n2)} = 3.50634892140253 * √((1/20) + (1/11)) = 1.316208338      

t = (x1-bar -x2-bar)/SE = 1.747443724       

p- value = 0.091143513       

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]